# Lattice QCD at finite temperature and density

### 17 Mar 2026 v1

@@ 2.3 Universality, scaling, and a generalized Banks–Casher relation
A concrete example is the improved order parameter $\tilde{M}$ proposed in recent work, schematically of the form $\tilde{M}=M_{\ell}-H\,\chi_{\ell}$, which suppresses the leading regular contribution in $H$ [81, 39, 80]. In this construction, universal scaling behavior becomes visible with controlled finite-volume effects. Such improved observables are particularly useful when using staggered ensembles at several pion masses down to $m_{\pi}\sim 55$ MeV.
@@ 3 Finite density: CEP status, phase boundary, and low-$T$ strategies
Figure 5: Status of lattice constraints and recent estimates for the CEP location. Left: exclusion region from Wuppertal–Budapest, including the bound of no CEP for $\mu_{B}<450~\mathrm{MeV}$ at $\sigma [22$]. Middle: Wuppertal–Budapest Lee–Yang-edge analysis, yielding an $84\%$ probability for no CEP below $T_{\rm CEP}<103~\mathrm{MeV}$ [2]. Right: Parma–Bielefeld Lee–Yang-edge extrapolation from imaginary-$\mu_{B}$ simulations [37].
@@
An emerging theme is to combine ideas from machine learning with field-theory algorithms to improve sampling in complex-weight systems. Diffusion-based generative models have been explored first in simple settings such as scalar field theories, and more recently in non-Abelian lattice gauge theories including two-dimensional SU(2) systems [91, 4]. In parallel, the worldvolume hybrid Monte Carlo algorithm provides a way to sample complexified manifolds related to Lefschetz thimbles. Both directions are under active development; their conceptual unification—controlled sampling on complexified manifolds combined with expressive probabilistic models—is promising for future applications [57, 58].
@@ 5 Thermodynamics in external conditions
At finite isospin chemical potential, QCD enters a pion-condensed BEC regime once the isospin density becomes energetically favored; in the convention commonly used in lattice studies, this occurs at $\mu_{I}\gtrsim m_{\pi}/2$, equivalently $\mu_{Q}\gtrsim m_{\pi}$ on the pure charge axis. Recent lattice determinations of the EoS at nonzero $\mu_{I}$ provide a controlled laboratory for finite-density physics, including characteristic structures of the interaction measure in the BEC phase and a stiffening of the EoS reflected in an enhanced speed of sound [27]. Going beyond the pure-$\mu_{I}$ axis, a leading-order Taylor expansion around simulation points at $\mu_{I}\neq 0$ enables access to small $\mu_{B}$ and $\mu_{S}$ deformations and, as a first application, yields the pressure along the pure charge axis $\mu_{Q}$, conveniently visualized as a three-dimensional surface $p/T^{4}=p/T^{4}(T,\mu_{Q})$ and compared to the direct $\mu_{I}$-axis results deep into the BEC regime [31, 27].
Two other externally controlled deformations are quark spin polarization and rotation. For a finite quark spin potential $\mu_{\Sigma}$, studied via analytic continuation from imaginary $\mu_{\Sigma}$, pseudocritical temperatures extracted from the Polyakov-loop susceptibility and from the disconnected chiral susceptibility follow a quadratic dependence on $\mu_{\Sigma}$; the resulting positive curvature implies that real spin polarization lowers both $T_{pc}^{\rm conf}$ and $T_{pc}^{\chi}$ [23]. For rotation, lattice simulations at imaginary angular velocity reveal a spatially inhomogeneous confinement–deconfinement structure in gluodynamics, with confined and deconfined regions separated by a radius-dependent boundary; analytic continuation to real rotation then suggests a deconfined core near the rotation axis and a confined periphery [24]. More recently, acceleration has also been explored and was likewise found to induce a spatially inhomogeneous confinement–deconfinement structure in gluodynamics, with confined and deconfined regions separated by a position-dependent boundary in Rindler spacetime; in this case, however, the boundary is approximately described by the Tolman–Ehrenfest law [25].

### 23 Mar 2026 v2

@@ 2.3 Universality, scaling, and a generalized Banks–Casher relation
A concrete example is the improved order parameter $\tilde{M}$ proposed in recent work, schematically of the form $\tilde{M}=M_{\ell}-H\,\chi_{\ell}$, which suppresses the leading regular contribution in $H$ [85, 41, 82, 83, 84]. In this construction, universal scaling behavior becomes visible with controlled finite-volume effects. Such improved observables are particularly useful when using staggered ensembles at several pion masses down to $m_{\pi}\sim 55$ MeV.
@@ 3 Finite density: CEP status, phase boundary, and low-$T$ strategies
Figure 5: Status of lattice constraints and recent estimates for the CEP location. Left: exclusion region from Wuppertal–Budapest, including the bound of no CEP for $\mu_{B}<450~\mathrm{MeV}$ at $2\sigma$ [23]. Middle: Wuppertal–Budapest Lee–Yang-edge analysis, yielding an $84\%$ probability for no CEP below $T_{\rm CEP}<103~\mathrm{MeV}$ [3]. Right: Parma–Bielefeld Lee–Yang-edge extrapolation from imaginary-$\mu_{B}$ simulations [39].
@@
An emerging theme is to combine ideas from machine learning with field-theory algorithms to improve sampling in complex-weight systems. Diffusion-based generative models have been explored first in simple settings such as scalar field theories, and more recently in non-Abelian lattice gauge theories including two-dimensional SU(2) systems [96, 1, 5]. In parallel, the worldvolume hybrid Monte Carlo algorithm provides a way to sample complexified manifolds related to Lefschetz thimbles. Both directions are under active development; their conceptual unification—controlled sampling on complexified manifolds combined with expressive probabilistic models—is promising for future applications [59, 60].
@@ 5 Thermodynamics in external conditions
At finite isospin chemical potential, QCD enters a pion-condensed BEC regime once the isospin density becomes energetically favored; in the convention commonly used in lattice studies, this occurs at $\mu_{I}\gtrsim m_{\pi}/2$, equivalently $\mu_{Q}\gtrsim m_{\pi}$ on the pure charge axis. Recent lattice determinations of the EoS at nonzero $\mu_{I}$ provide a controlled laboratory for finite-density physics, including characteristic structures of the interaction measure in the BEC phase and a stiffening of the EoS reflected in an enhanced speed of sound [28]. Going beyond the pure-$\mu_{I}$ axis, a leading-order Taylor expansion around simulation points at $\mu_{I}\neq 0$ enables access to small $\mu_{B}$ and $\mu_{S}$ deformations and, as a first application, yields the pressure along the pure charge axis $\mu_{Q}$, conveniently visualized as a three-dimensional surface $p/T^{4}=p/T^{4}(T,\mu_{Q})$ and compared to the direct $\mu_{I}$-axis results deep into the BEC regime [32, 28]. Complementary insight into the onset of pion condensation at finite isospin chemical potential has also been obtained from an unbiased exponential resummation of Lee–Yang zeros, yielding estimates compatible within uncertainties with the pion-condensation threshold [86].
Two other externally controlled deformations are quark spin polarization and rotation. For a finite quark spin potential $\mu_{\Sigma}$, studied via analytic continuation from imaginary $\mu_{\Sigma}$, pseudocritical temperatures extracted from the Polyakov-loop susceptibility and from the disconnected chiral susceptibility follow a quadratic dependence on $\mu_{\Sigma}$; the resulting positive curvature implies that real spin polarization lowers both $T_{pc}^{\rm conf}$ and $T_{pc}^{\chi}$ [24]. For rotation, lattice simulations at imaginary angular velocity reveal a spatially inhomogeneous confinement–deconfinement structure in gluodynamics, with confined and deconfined regions separated by a radius-dependent boundary; analytic continuation to real rotation then suggests a deconfined core near the rotation axis and a confined periphery [25]. Recently, acceleration was first studied in lattice gluodynamics and found to strongly soften the deconfinement transition under weak acceleration [35]; this was later reformulated in Rindler spacetime, yielding qualitatively similar behavior together with a spatially inhomogeneous confinement–deconfinement structure, whose boundary approximately follows the Tolman–Ehrenfest law [26].